OpenStudy (nuccioreggie):

A set of equations is given below: Equation C: y = 4x + 8 Equation D: y = 4x + 2 Which of the following best describes the solution to the given set of equations? No solution One solution Two solutions Infinitely many solutions

OpenStudy (nuccioreggie):

@jameshorton

OpenStudy (jameshorton):

c

OpenStudy (jchick):

y=x+14 line 1 y=3x+2 line 2 These are both the equation of lines written in slope intercept form y=mx+b where m is the slope and the point (0,b) is the y intercept. The first line has a slope of m=1. The 2nd line has a slope of m=3 Since these lines have different slopes, they are not parallel, thus they will cross at some point. What you have to determine is where the lines cross, which will be a point (x,y) that is on both lines. We already have y solved in terms of x from either equation so we can use substitution to solve the system. Since y=x+14 from line 1, put x+14 in place of y in the equation of line 2. x+14=3x+2 solve for x. Subtract x from both sides... 14= 3x-x+2 14=2x+2 subtract 2 from both sides 14-2=2x 12=2x divide both sides by 2 6=x We now have the x value of the common point. Plug the value 6 in for x in one of the original equations and solve for y. y=6+14 y=20 These two lines cross at the point (6,20) which is a point the two lines have in common.

OpenStudy (jchick):

This is for your last one.

OpenStudy (jameshorton):

huh what you mean

OpenStudy (jchick):

He posted a different question a second ago and said he thought it was A.

OpenStudy (jameshorton):

oh when

OpenStudy (jchick):

Just a few moments ago and you said he was right.

OpenStudy (jchick):

I just replied to it so you should be able to see that I did.